OpenStudy (anonymous):

is y=0 a horizontal line

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Y=0 is a point on a graph, in order to tell if it's a horizontal line, you need a second point.

OpenStudy (anonymous):

how about 3x+2y=o

OpenStudy (mathstudent55):

y = mx + b is the equation of a line with m slope and (0, b) as the y-intercept. y = 0 can be thought of y = 0x + 0, with m = 0, and b = 0, so it's a line with slope = 0 that passes though (0, 0).

OpenStudy (mathstudent55):

3x + 2y = 0 can be solved for y and put in the y = mx + b form: 3x + 2y = 0 2y = -3x y = (-3/2)x Since y = (-3/2)x = (-3/2)x + 0, m = -3/2 and y-intercept (0, 0)